LARGE PARITY SPLITTING AS AN INDICATION OF THE HCP ⟷ CPH “ISOMERIZATION”

\underline{HARUKI ISHIKAWA}, Department of Molecular Science and Material Engineering, Graduate School of Science and Technology, Kobe University, Kobe 657-8501, Japan; YASUHIKO MURAMOTO, and NAOHIKO MIKAMI, Department of Chemistry, Graduate School of Science, Tohoku University, Sendai 980-8578, Japan.

We have been investigating the HCP ⟷ CPH “isomerization” of HCP molecule by the stimulated emission pumping (SEP) spectroscopy. In the previous study, we recorded SEP spectra in the energy region up to 26800 cm^{−1} in the X̃ state from the vibrational ground level. This value is very close to the height of the “isomerization” barrier (27400 cm^{−1}) predicted by the theoretical study. It was found that several vibrational levels exhibit complicated perturbation patterns due to the Coriolis type interaction^{a}. Strong Coriolis perturbations in the energy region near the “isomerization” barrier are predicted by Jacobson and Child based on their spherical pendulum model^{b}. They also pointed out that the parity splitting in ℓ ≠ 0 levels should be a good spectroscopic indication of the “isomerization” in addition to the rotational constants. Until now, due to the rotational selection rules, only the ℓ = 0 levels were observed in the energy region above 20000 cm^{−1}. However, using a perturbed level in the C̃ state as an intermediate, we have succeeded in recording SEP spectra which have sampled ℓ = 1 levels in the 26250 – 27300 cm^{−1} region in the X̃ state, in the present study. As a result, we have found a good relation between the rotational constants and the parity splittings. Larger-B levels exhibit larger parity splittings. For instance, A vibrational level having a B_{ave.} of 0.76 cm^{−1} exhibits ΔB = 0.13cm^{−1}, while another level having a B_{ave.} of 0.62 cm^{−1} exhibits ΔB = 0.03cm^{−1}. The ΔB of the former level is about 4 times larger than that of the latter. It is confirmed that the parity splitting is one of the spectroscopic indications of the “isomerization” of the HCP/CPH system as expected.

aY. Muramoto, H. Ishikawa, and N. Mikami, 60th Ohio State University International Symposium on Molecular Spectroscopy, WH03 (2005).

bM. P. Jacobson and M. S. Child, J. Chem. Phys. \underline{114}, 262 (2001).